OpenStudy (anonymous):

find the solutions... tan x + sec x = 2cosx x[0,2pi]

OpenStudy (anonymous):

@Mani_Jha @Ishaan94 @Zarkon @experimentX

OpenStudy (anonymous):

sinx + 1 = 2 - 2sin^2x

OpenStudy (anonymous):

Solve it now.

OpenStudy (callisto):

tan x + sec x = 2cosx (sinx +1) / cosx = 2cosx sinx +1 = 2cos^2 x sinx +1 = 2(1-sin^2x) 2sin^2x +sinx -1 =0 ...

OpenStudy (anonymous):

i am getting sin x = 1/2 and -1....is it right????

OpenStudy (anonymous):

It is

OpenStudy (anonymous):

ohhh i forgot...there is a range of the function....the full question has been modified above

OpenStudy (anonymous):

sin3pi/2 = -1 sinpi/6 = 1/2

OpenStudy (mani_jha):

you could also multiply L.H.S. and R.H.S with (secx-tanx)

OpenStudy (anonymous):

hmm 3solutions